OpenStudy (anonymous):

Suppose that the cost C(x), in dollars, of producing x compact discs is given by C(x)=0.5x+500. A) graph B) Compute C(150) C)Compute the cost of the next compact disc.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I am just not sure how to to put this into the calc to do the graph

OpenStudy (anonymous):

what kind of calc do you have - i have ti 84 plus

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

on top line (directly below screen) hit Y=

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

enter in 0.5x +500 then hit graph key

OpenStudy (anonymous):

i went to website for some cool downloads and manual

OpenStudy (anonymous):

OK I need to do that

OpenStudy (anonymous):

cool stuff works well with chemistry also :)

OpenStudy (anonymous):

what site did you use

OpenStudy (anonymous):

it is not graphing anything

OpenStudy (anonymous):

These are the x and y intercept of the equation y=0.5x +500 y=500,1x=-1000

OpenStudy (anonymous):

here is another good calculator site http://www.mathway.com

OpenStudy (anonymous):

hit zoom

OpenStudy (anonymous):

how did you get -1000

OpenStudy (anonymous):

that worked

OpenStudy (anonymous):

set y = 0 and solve for x

OpenStudy (anonymous):

So I divide 500 by 0.5

OpenStudy (anonymous):

To find the x-intercept substitute in 0 for y and solve for x. (0)=0.5x+500 Since x is on the right-hand side of the equation switch the sides so it is on the left-hand side of the equation 0.5x+500=0 Since 500 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 500 from both sides. 0.5x=-500 Divide each term in the equation by 0.5. 0.5x/0.5=-500/0.5 Cancel the common factor of 0.5 1x=-500/0.5 Divide -500 by 0.5 to get -1000. 1x=-1000 To find the y-intercept, substitute in 0 for x and solve for y. y=0.5(0)+500 Multiply 0.5 by each term, inside the parentheses y=0+500 Solve the equation y=500

OpenStudy (anonymous):

this will give you the answer I gave you above These are the x and y intercept of the equation y=0.5x +500 y=500,1x=-1000

OpenStudy (anonymous):

This is what I got. I just did it different. Is there other ways to do this?

OpenStudy (anonymous):

Thank you for your time.

OpenStudy (anonymous):

not really sure if there are other ways but you are welcome for the time

OpenStudy (anonymous):

For the next part do I just plug the C(150) the 150 in place of the x?

OpenStudy (anonymous):

yes 150 in place of x